Hey there! I would like to know how to do problems like how to find the length of the line segment whose end points are (6,7) and (8,0). I would really appreciate it if I you would provide an explanation (no hard terms please!) on how you would answer my problem. I've been to so many sites but I don't really get how they explained my topic because they provided really hard-to-get explanations with complex terms and I was hoping you would be able to help me understand the problem. Thank you and I am very much looking forward to your reply. :D
Answer by stanbon(75887)

find the length of the line segment whose end points are (6,7) and (8,0).
--
Formula: distance = sqrt[(x1-x2)^2+(y1-y2)^2]
------
Your Problem:
d = sqrt[(6-8)^2+(7-0)^2] = sqrt[2^2 + 7^2] = sqrt(53)
===============
Cheers,
Stan H.
========